OpenStudy (anonymous):

Solve the following equations \[X ^{2} +4x-11=0\]

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ x^2+4x+11=0 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

what do you mean@mathmath333

OpenStudy (mathmath333):

i mean is the \(x\) in small case letters

OpenStudy (mathmath333):

by squaring method \(\large \begin{align} \color{black}{ x^2+4x+11=0 \hspace{.33em}\\~\\ \implies x^2+2\cdot 2x+2^2-2^2+11=0 \hspace{.33em}\\~\\ \implies x^2+4x+4-4+11=0 \hspace{.33em}\\~\\ \implies (x+2)^2+7=0 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

see if it makes sense so far

OpenStudy (anonymous):

thx

OpenStudy (mathmath333):

lol its not the full answer its only partially solved

OpenStudy (mathmath333):

continuing further \(\large \begin{align} \color{black}{ (x+2)^2+7=0 \hspace{.33em}\\~\\ (x+2)^2-(-7)=0 \hspace{.33em}\\~\\ (x+2)^2-(\sqrt{7i})^2=0 \hspace{.33em}\\~\\ \left( x+2-\sqrt{7}i\right) \times (x+2+\sqrt{7}i)=0 \hspace{.33em}\\~\\ x=-2+\sqrt{7}i~~~or~~~x=-2-\sqrt{7}i \hspace{.33em}\\~\\ }\end{align}\)

Nnesha (nnesha):

lol mathmat3333